OpenStudy (thecalchater):

Can anyone help with a few calc questions? A rectangular pare is to contain 81 square inches of print. The margins on each side are 1 inch. Find the dimentiond of the page such taht the least amount of paper is used. A. 11,11 B.12,12 C. 13,13 D. 10,10 Find the differential of the function y=xcos(8x) A. cos(8x)dx B.-8xsin(8x)dx C.cos(8x)-8xsin(8x)dx D.(cos(8x)+8xsin(8x)dx

OpenStudy (thecalchater):

Also can you check a few. Use Newton's Method to approximate the zero of the function f(x)=x^3+2x+2 I got -.7709 The measurement of the edge of a cube is found to be 5 inches with a possible error of .005 inch. Use differentials to estimate the propagated error in computing a the volume and b the surface area. I got .375 cubic inches and .3 square inches

OpenStudy (thecalchater):

@Loser66

OpenStudy (utterly_confuzzled):

The second question would be y=xcos(8x) y'=x(-sin(8x)(8))+ cos(8x) y'= cos(8x) - 8xsin(8x)

OpenStudy (thecalchater):

ok thanks. and the first one? I am thinking B

OpenStudy (utterly_confuzzled):

And I would assume that the first on is D, simply because 9x9=81, so if the page has 1inch margins, and 10-1=9, then wouldn't that use the least amount of paper??

OpenStudy (thecalchater):

can you check my other 2?

OpenStudy (utterly_confuzzled):

 I did the math wrong orginaly, it would be a, because you lose 2 inches from the total width, not 1